OpenStudy (fairy733189):

A right triangle has base (x - 5) units and height (4x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. Part B: What is the area of the triangle? Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials.

OpenStudy (wilsondanielle):

Because it's a right triangle, you can use the Pythagorean theorem - a^2 + b^2 = c^2. You have a (x-5) and b (4x-10) - from there, you need to find c (for part a)

OpenStudy (fairy733189):

so I multiply a and b to get c to solve part A?

OpenStudy (wilsondanielle):

a squared and b squared multiplied will give you c squared - simply a times b will not give you a correct answer.

OpenStudy (fairy733189):

so for part a what would i need to do?

OpenStudy (wilsondanielle):

you need the product of a^2 and b^2 - so find that, and you'll have c^2. From there, how do you find c?

OpenStudy (fairy733189):

im not really sure

OpenStudy (fairy733189):

can you explain how?

OpenStudy (wilsondanielle):

another way to phrase the theorem is a*a + b*b = c*c . Does that make it clear?

OpenStudy (fairy733189):

yes. I understand the pythagorean theorem but what I don't understand is when I get c^2, is that my answer for part a?

OpenStudy (wilsondanielle):

no, you need to take the square root of c^2 to get c, and that will be your answer for part a.

OpenStudy (mathmale):

"A right triangle has base (x - 5) units and height (4x - 10) units." If you square the two given dimensions (which represent the lengths of the two sides of the triangle), and then add them together, the result is the answer you want. Please review the original question.

OpenStudy (mathmale):

"What is the square of the length of the hypotenuse of the triangle? " You want c^2, not c.

OpenStudy (mathmale):

So you had the cprrect answer all along.

OpenStudy (fairy733189):

So this is how i would do it: a^2 + b^2 = c^2 (x-5)^2 + (4x-10)^2 x^2 - 10 + 25 + 16x^2 - 80x + 100

OpenStudy (fairy733189):

Is this right so far?

OpenStudy (mathmale):

Thanks for sharing your work. Wouldn't that 2nd term be -10x, not just -10? Otherwise you're right on track.

OpenStudy (fairy733189):

your right, sorry I didn't realize that. ANd after that I add them to get c^2?

OpenStudy (mathmale):

Combine like terms, and then you'll be done with Part A.

OpenStudy (mathmale):

Combine would be a better term here than Add.

OpenStudy (mathmale):

Or, Simplify.

OpenStudy (mathmale):

Part B: You are given both the base and the height of this triangle. What's the formula for the area of a triangle, when such info is given?

OpenStudy (fairy733189):

x^2 - 10x + 25 + 16x^2 - 80x + 100 17x^2 - 90x +125 Is this right?

OpenStudy (fairy733189):

the formula would be base times height so I would multiply (x - 5)(4x - 10)?